# Implicit Gaussian process representation of vector fields over arbitrary latent manifolds

### v1

@@ 4.2 Scalable training via inducing point methods
A notable drawback of GPs is their computational inefficiency, largely driven by the need to compute an $n\times n$ covariance matrix. Therefore, inducing point methods, which reduce the effective number of data points to a set of $n^{\prime}<n$ inducing points, have become a mainstay of GPs in practice (Titsias, 2009). Hutchinson et al. (2021) showed that the inducing point methods such as that of Titsias (2009) are readily extendible to vector fields over Riemannian manifolds, provided the covariance matrix of inducing points is represented in tangent space coordinates. Since the kernel of RVGP is constructed from a positional encoding using connection Laplacian eigenvectors, which are expressed in local coordinates, our method is readily amenable to computation by inducing point methods which we provide an implementation of.
@@ 5.1 Manifold-consistent interpolation of vector fields
Figure 2: Superresolution and inpainting. A Sparse samples randomly distributed over the Stanford bunny and torus surface mesh are interpolated to a higher resolution ($k=50$). B Ablation studies for the Stanford Bunny, showing the dependence of mean alignment of the superresolved vector fields in the test set against the average distance between manifold points $\alpha$ (for $k=50$ fixed) and the number of eigenvectors $k$. C Prediction in a masked area containing vector field singularity. Predictions using an RBF kernel with Laplacian eigenvectors as positional encoding produce discontinuities along the masked boundary (triangle), vectors that protrude the mesh surface (star) and do not converge to zero magnitude at the singularity. RVGP using connection Laplacian eigenvectors produces smoothly varying inpainting ($k=50$).
We expected that RVGP fitted to sparse vector samples over an unknown manifold would leverage the global regularity of the vector field to provide accurate out-of-sample predictions. Thus, we conducted two experiments on the Stanford Bunny and toroidal surface mesh to test the RVGP’s ability to super-resolve sparse samples and inpaint missing regions containing singularities on diverse manifold topologies. In both experiments, we first generated a ground truth vector signal. Next, we drew $n$ evenly sampled anchor points $\{({\bm{x}}_{i}\})$ using the furthest point sampling algorithm (Qi et al., 2017), which has a parameter $\alpha$ quantifying the average pairwise distance of points relative to the manifold diameter. We then sampled corresponding vectors $\{{\bm{v}}_{i}\}$ from a uniform distribution on the sphere $S^{3}$, projected them onto their respective tangent spaces $\hat{{\bm{v}}}_{i}={\mathbb{T}}_{i}^{T}{\bm{v}}_{i}$ and used the vector heat method (Sharp et al., 2019) to find the smoothest vector field. Specifically, concatenating signals as $\hat{{\bm{v}}}=\mathbin{\|}_{i=0}^{n}\hat{{\bm{v}}}_{i}\in\mathbb{R}^{nd\times 1}$ the vector heat method obtains $\hat{{\bm{v}}}(\tau)=\hat{{\bm{v}}}(\tau)/u(\tau)$, where $\hat{{\bm{v}}}(\tau)$ is the solution of the vector heat equation $\hat{{\bm{v}}}(\tau)=\hat{{\bm{v}}}\exp{\left(-{\bm{L}}_{c}\tau\right)}$ and $u(\tau)$ is the solution of the scalar heat equation $u(\tau)=|\hat{{\bm{v}}}|\exp{\left(-{\bm{L}}\tau\right)}$. To obtain a smooth field, we ran the process until diffusion time $\tau=100$. On surfaces of genus 00 this process will lead to at least one singularity.
@@
In the first experiment, we asked if RVGP can smoothly interpolate (also known as super-resolve) the vector field from sparse samples. To this end, we fitted a graph over the bunny mesh and trained RVGP using vectors over $50\%$ of the nodes, reserving the remaining $50\%$ for testing. We found that RVGP predictions were in excellent visual alignment with the training vector field (Fig. 2A). To investigate the sensitivity of predictions to sampling density and manifold description, we conducted ablation studies by varying both $\alpha$ and the number of eigenvectors $k$. As quantified by the inner product between predicted and test vectors, we found that RVGP produces accurate alignment with only $10$ eigenvectors over a broad range of sampling densities (Fig. 2B).
@@
In the second experiment, we tested RVGP’s ability to inpaint whole vector field regions containing singularities. This experiment is more challenging than classical inpainting because it requires our method to infer the smoothest topologically consistent vector field. To this end, we masked off the vortex singularity and used the remaining points to train RVGP consisting of a positional encoding of the manifold based on graph Laplacian eigenvectors. For comparison, we also trained a radial basis function kernel that treats vector entries independently. We found that vectors predicted by RVGP closely followed the mesh surface, aligned with the training set on the mask boundary and smoothly resolved the singularity by gradually reducing the vector amplitudes to zero. By contrast, the RBF kernel did not produce vectors with smooth transitions on the mask boundary and often protruded from the mesh surface. This example thus shows that the connection Laplacian positional encoding with vector field kernel provides sufficient regularity to learn vector fields over complex shapes.
@@ 5.2 Superresolution of EEG data
Finally, as a biologically and clinically relevant use case, we applied RVGP to superresolve electroencephalography (EEG) recordings from human subjects. EEG recordings measure spatiotemporal wave patterns associated with neural dynamics, which play a fundamental role in human behaviour (Sato et al., 2012; Xu et al., 2023). Accurately resolving these dynamics requires high-density EEG setups in excess of 200 channels Robinson et al. (2017); Seeber et al. (2019); Siclari et al. (2018). However, due to long setup times during which signal quality can rapidly degrade, experimentalists and clinicians commonly resort to low-density recordings with 32 or 64 channels (Chu, 2015).
@@
Thus, we asked whether superresolving low-density 64-channel recordings (Fig. 3A) using RVGP into high-density 256-channel recordings (Fig. 3C) can facilitate biological discovery and clinical diagnostics. As a ground truth, we collected 256-channel EEG recordings of resting-state brain activity from 33 Alzheimer’s patients (AD) and 28 age-matched healthy controls (see Appendix A.1). We focused on low-frequency alpha waves spanning the 8-15 Hz range, which represent the dominant rhythm at rest (Berger, 1934) and exhibits impaired dynamics in AD (Moretti et al., 2004; Besthorn et al., 1994; Dauwels et al., 2010). After preprocessing of EEG time series (see Appendix A.1), we used a triangulated mesh of the head surface, with vertices given by 256 electrode locations, and finite differencing to compute the vector field at mesh vertices (Fig. 3B,C (ground truth), Appendix A.2). Then, we constructed RVGP kernels using the connection Laplacian eigenvectors derived from a $k$-nearest neighbour graph ($k=5$) fit to mesh vertices. Finally, we trained RVGP using vectors at 64 training vertices (Fig. 3B) and then used it to infer the vector field at the remaining 192 test vertices (Fig. 3C). As a benchmark, we also used a channel-wise linear and spline interpolation of the vector field.
@@
Figure 3: Reconstruction of spatiotemporal wave patterns in the human EEG. A Snapshot of an alpha wave pattern (8-15 Hz) recorded on low density (64 channel) EEG from a healthy subject projected in two dimensions. B Phase field of an alpha wave. Vector field denotes the the spatial gradient of the voltage signal. C Ground-truth and reconstructed high-density phase field (256 channel) using RVGP, linear and spline interpolation. Streamlines, computed based on the vector field, highlight features of the phase field. RVGP significantly better preserves singularities, i.e., sources, sinks and vortices. D Reconstruction accuracy, measured by the preservation of singularities. E Receiver operating characteristic (ROC) for binary classification of patients with Alzheimer’s disease against healthy controls using a linear support vector machine trained on the divergence and vorticity fields.
@@
To assess the quality of reconstructions, we computed the divergence and curl of the predicted EEG vector field and computed the mean absolute error (MAE) relative to the ground truth 256-channel EEG. The divergence and curl have previously been used to identify singularities in neural wave patterns from human neuroimaging recordings and have been linked to cognitive function and behaviour (Roberts et al., 2019; Xu et al., 2023). We found that the RVGP reconstruction closely approximated the divergence and curl of the high-density EEG (Fig. 3C). For visual comparison, we show a single time snapshot with characteristic vector field singularities such as sources, sinks and vortices. In particular, RVGP reconstruction substantially better preserved these singularities than linear interpolation, which had the tendency to underfit the vector field, and spline interpolation, which introduced spurious local structures. This observation is corroborated by significantly lower angular error and curl-divergence MAE for all subjects for RVGP compared with benchmark ($n=61$, Fig. 3D).
@@
Given the superior reconstruction accuracy of our approach, we sought to determine if it could enhance the classification accuracy between patients with Alzheimer’s Disease (AD) and healthy controls. Contemporary diagnostic methods for AD are costly, invasive, and lengthy in procedure (Zetterberg & Bendlin, 2021). To underscore the potential clinical significance and scalability of our methodology, we employed a linear support vector machine for the classification of AD patients versus age-matched controls. This classification was based on the reconstructed divergence and curl fields derived from a brief, one-minute resting-state low-density EEG recording – a procedure that is feasibly integrated into clinical settings due to its non-invasive nature and cost-efficiency. Our results show significantly higher than state-of-the-art classification accuracy, approaching that derived from the ground truth high-density EEG signal (Figure 3E).
6 Conclusion
@@
In this study, we introduced RVGP, a novel extension of Gaussian processes designed to model vector fields on latent Riemannian manifolds. Utilising the spectrum of the connection Laplacian operator, RVGP intrinsically captures the manifold’s geometry and topology and the vector field’s smoothness. This enables the method to learn global patterns while preserving singularities, thereby filling a significant gap in existing approaches that are limited to known, analytically tractable manifolds. A key strength of RVGP is its data-driven, intrinsic construction via a proximity graph, which enhances its practical utility by making it highly applicable to real-world datasets where the manifold is often not explicitly known. Demonstrated across diverse scientific domains such as neuroscience and geometric data analysis, RVGP not only advances the field of geometrically-informed probabilistic modelling but also offers a statistical tool for various high-impact applications, including clinical neuroscience, that is robust to sampling density and noisy data.
Author Contributions
AG and PV devised the RVGP method and developed the theoretical foundations with contributions from RLP. AG and RLP wrote the RVGP code. AG, RLP and MVC designed the experiments. MVC and NG conceptualised the EEG preprocessing pipeline. MD, EM, DS, and PAM collected the EEG data. MVC, RLP and NG performed the analysis of the EEG data with contributions from AG. AG, RLP, MVC and PV wrote the paper.
Acknowledgments
We thank Nicolas Aspert for the critical reading and feedback on the manuscript. AG acknowledges support from an HFSP Cross-disciplinary Postdoctoral Fellowship (LT000669/2020-C). PAM acknowledges funding from the NIHR Biomedical Research Centre at Imperial College London. RLP acknowledges the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project-ID 424778381-TRR 295.
Code availability
The Python code to reproduce our results is openly available at https://github.com/agosztolai/RVGP.
@@ A.1 Experimental data, and pre-processing
Experimental data: We analysed 10-minute resting state EEG recordings from a cohort of 61 subjects containing 33 patients with a clinical diagnosis of AD and 28 age-matched healthy control subjects. Data was acquired from the clinical imaging facility at Hammersmith Hospital in London. The AD group consisted of 15 Females and 18 Males with a mean age of 76 years (std = 7.8 years), healthy control group contained 15 Females and 13 Males with a mean age of 77 (std = 4.9 years). 10 participants were excluded from the original full cohort of 71 subjects due to poor data quality. During pre-processing, the EEG was bandpass filtered from 2-40 Hz with a second-order Butterworth filter and downsampled to 250Hz. Channels exceeding a kurtosis threshold of 3 were rejected before average re-referencing. We performed an independent component analysis with the Picard algorithm and applied an automatic artefact rejection algorithm in eeglab with a maximum of 5 % of components removed from each dataset. The pre-processed EEG was then filtered in the alpha band (8-15Hz). The low-density EEG consisted of 64 channels sampled evenly across the scalp. These channels were then used to reconstruct the full 256-channel high-density recording.
@@ A.4 Feature Extraction
To determine the behaviourally relevant regions of the cortex, we computed the time-resolved divergence and vorticity fields for each subject. Next, we estimated the probability density of sources, sinks and spirals by computing the mean positive/negative vorticity and mean positive/negative divergence fields thresholded at $1$ and $<-1$ averaged over time for each node. This produced four cortical probability maps per subject; mean source probability (positive divergence, $D>1$), mean sink probability (negative divergence $D<-1$), clockwise spiral probability (positive vorticity, $C>1$), anticlockwise spiral probability (negative vorticity, $C<-1$). We extracted two predictors from these four probability maps. The first predictor was the probability ratio of sources to sinks at a given node. The second predictor was the probability ratio of clockwise to anti-clockwise spirals at a given node:
@@
where $P_{n}(A)$ represent either the probability of a source or a clockwise spiral at node $n$ and $P_{n}(B)$ represents the probability of a sink or an anti-clockwise at node $n$. We ran a group-level cluster permutation analysis to identify nodes where the source-to-sink ratio or clockwise-to-anti-clockwise spiral ratio differed between the patients with AD and healthy controls in the reference EEG (alpha = 0.05, 50000 permutation). The three clusters with the lowest p-value were then used as seed regions where we extracted the source-to-sink probability ratio and clockwise to anti-clockwise spiral probability ratio in the interpolated EEG for each reconstruction method (RVGP, spline and linear interpolation).

### v2

@@ 4.2 Scalable training via inducing point methods
A drawback of GPs is their computational inefficiency, driven by the need to compute an $n\times n$ covariance matrix. Therefore, inducing point methods, which reduce the effective number of data points to a set of $n^{\prime}<n$ inducing points, have become a mainstay of GPs in practice (Titsias, 2009). Hutchinson et al. (2021) showed that inducing point methods, e.g., Titsias (2009), are extendible to vector fields over Riemannian manifolds, provided the covariance matrix of inducing points is represented in tangent space coordinates. Since the kernel of RVGP is constructed from a positional encoding using connection Laplacian eigenvectors, which are expressed in local coordinates, our method is readily compatible with inducing point methods, which we provide an implementation of.
@@ 5.1 Manifold-consistent interpolation of vector fields
Figure 2: Superresolution and inpainting. A Matrix-valued kernel (RVGP) against a scalar-valued kernel (e.g., in Hutchinson et al. (2021)). B Uniformly distributed samples over the Stanford bunny and torus are interpolated to a higher resolution ($k=50$). C Ablation studies for the Stanford bunny, showing the dependence of alignment of the superresolved vector fields in the test set against data density quantified by the average distance between manifold points $\alpha$ (for $k=50$ fixed) and the number of eigenvectors $k$. The vectorial RVGP representation is compared against a channel-wise representation using an RBF kernel with Laplacian eigenvectors as positional encoding. D Prediction of singularity in masked area. RBF kernel predicts discontinuities along the masked boundary (triangle), vectors that protrude the mesh surface (star) and do not converge to zero magnitude at the singularity. RVGP predicts smoothly varying inpainting ($k=50$).
We expected that RVGP fitted to sparse vector samples over an unknown manifold would leverage the global regularity of the vector field to provide accurate out-of-sample predictions. Thus, we conducted two experiments on the Stanford bunny and toroidal surface mesh to test the RVGP’s ability to super-resolve sparse samples and inpaint missing regions containing singularities on diverse manifold topologies. Given $n$ uniformly sampled anchor points $\{{\bm{x}}_{i}\}$ on the surface mesh, we generated a smooth ground truth vector signal over these points by sampling vectors $\{{\bm{v}}_{i}\}$ from a uniform distribution on the sphere $S^{2}$, projecting them onto their respective tangent spaces $\hat{{\bm{v}}}_{i}={\mathbb{T}}_{i}^{T}{\bm{v}}_{i}$ and using the vector heat method (Sharp et al., 2019) to find the smoothest vector field. Specifically, concatenating signals as $\hat{{\bm{v}}}=\mathbin{\|}_{i=0}^{n}\hat{{\bm{v}}}_{i}\in\mathbb{R}^{nm\times 1}$ the vector heat method obtains $\hat{{\bm{v}}}\mapsto\hat{{\bm{v}}}(\tau)/u(\tau)$, where $\hat{{\bm{v}}}(\tau)=\hat{{\bm{v}}}\exp{\left(-{\bm{L}}_{c}\tau\right)}$ is the solution of the vector heat equation and $u(\tau)=|\hat{{\bm{v}}}|\exp{\left(-{\bm{L}}\tau\right)}$ is the solution of the scalar heat equation. We ran the process until diffusion time $\tau=100$. On surfaces of genus 00 this process will lead to at least one singularity.
@@
First, we asked if RVGP can smoothly interpolate (also known as super-resolve) the vector field from sparse samples. To this end, we fitted a graph over the bunny mesh and trained RVGP using vectors over $50\%$ of the nodes, holding out the rest for testing. For benchmarking, we also trained a radial basis function (RBF) kernel that treats vector entries as independent scalar channels. We found that RVGP predictions were in excellent visual alignment with the training vector field for dense (Fig. 2B) and sparse data (Fig. S2). The predictions of our model remain accurate when the points do not lie on the manifold surface but are drawn from a distribution centred on the manifold. Indeed, when we added Gaussian geometric noise of increasing standard deviation to manifold points, we found that the prediction accuracy was largely unaffected (Fig. S3).
We conducted two ablation studies to further investigate the robustness of RVGP predictions. First, we subsampled the surface mesh using the furthest point sampling algorithm (Qi et al., 2017) to simultaneously reduce the resolution and the amount of data used for training. Here, a parameter $\alpha$ controls the average pairwise distance of points relative to the manifold’s diameter. As quantified by the inner product between predicted and test vectors, RVGP produced accurate alignment with only $10$ eigenvectors over a broad range of sampling densities (Fig. 2C). By contrast, the benchmark method suffered a drop in performance at lower densities (higher $\alpha$). Next, we found that on a high-resolution surface ($\alpha=1.5\%$), RVGP yields high accuracy already with a few eigenvectors ($k$) which progressively increased with $k$. By contrast, the benchmark achieved good representation only from $k=50$ eigenvectors, thus achieving inferior dimensionality reduction.
@@
In the second experiment, we tested RVGP’s ability to inpaint whole vector field regions containing singularities. This experiment is more challenging than classical inpainting because it requires our method to infer the smoothest topologically consistent vector field. To this end, we masked off the vortex singularity and used the remaining points to train RVGP. We found that vectors predicted by RVGP closely followed the mesh surface, aligned with the training set on the mask boundary and smoothly resolved the singularity by gradually reducing the vector amplitudes to zero (Fig. 2D). By contrast, the RBF kernel did not produce vectors with smooth transitions on the mask boundary and often protruded from the mesh surface, showing that vectors treated as independent scalar fields do not capture the geometry of the tangent spaces. Thus, the connection Laplacian positional encoding provides sufficient regularity to learn vector fields over complex shapes.
@@ 5.2 Superresolution of EEG data
Finally, as a biologically and clinically relevant use case, we applied RVGP to superresolve electroencephalography (EEG) recordings from humans. EEG recordings measure spatiotemporal wave patterns associated with neural dynamics, which play a fundamental role in human behaviour (Sato et al., 2012; Xu et al., 2023). Accurately resolving these dynamics requires high-density EEG setups in excess of 200 channels (Robinson et al., 2017; Seeber et al., 2019; Siclari et al., 2018). However, due to long setup times during which signal quality can rapidly degrade, experimentalists and clinicians commonly resort to low-density recordings with 32 or 64 channels (Chu, 2015).
@@
Thus, we asked whether superresolving low-density 64-channel recordings (Fig. 3A) using RVGP (Fig. 3C) can facilitate biological discovery and clinical diagnostics. As a ground truth, we collected 256-channel EEG recordings of resting-state brain activity from 33 Alzheimer’s patients (AD) and 28 age-matched healthy controls (see Appendix A.1). We focused on low-frequency alpha waves spanning 8-15 Hz, which represent the dominant rhythm at rest (Berger, 1934) and exhibits impaired dynamics in AD (Moretti et al., 2004; Besthorn et al., 1994; Dauwels et al., 2010). After preprocessing of EEG time series (see Appendix A.1), we used a triangulated mesh of the scalp, with 256 known electrode locations as vertices, and finite differencing to compute the corresponding vectors (Fig. 3B,C (ground truth), Appendix A.2). Then, we constructed RVGP kernels using the connection Laplacian eigenvectors derived from a $K$-nearest neighbour graph ($K=5$) fit to vertices. Finally, we trained RVGP using vectors at 64 training vertices (Fig. 3B) and used it to infer the vectors at the remaining 192 test vertices (Fig. 3C). As a benchmark, we used a channel-wise interpolation of the vector field using linear, spline and RBF kernel methods.
@@
Figure 3: Reconstruction of spatiotemporal wave patterns in human EEG. A Snapshot of an alpha wave pattern (8-15 Hz) recorded on low density (64 channel) EEG from a healthy subject projected in two dimensions. B Phase field of an alpha wave. Vector field denotes the the spatial gradient of the voltage signal. C Ground-truth and reconstructed high-density phase field (256 channel) using RVGP, linear and spline interpolation. Streamlines, computed based on the vector field, highlight features of the phase field. RVGP significantly better preserves singularities, i.e., sources, sinks and vortices. D Reconstruction accuracy, measured by the preservation of singularities. E Receiver operating characteristic (ROC) for binary classification of patients with Alzheimer’s disease against healthy controls using a linear support vector machine trained on the divergence and vorticity fields. Shaded areas indicate a 95% confidence interval.
@@
To assess the quality of reconstructions, we computed the divergence and curl of the predicted EEG vector field and computed the mean absolute error (MAE) relative to the ground truth 256-channel EEG. Divergence and curl have previously been used to identify singularities in neural wave patterns in human neuroimaging and have been linked to cognitive function and behaviour (Roberts et al., 2019; Xu et al., 2023). We found that the RVGP reconstruction closely approximated the divergence and curl of the high-density EEG (Fig. 3C). For visual comparison, we show a snapshot with characteristic vector field singularities such as sources, sinks and vortices, which were significantly better preserved by RVGP than benchmarks. Linear interpolation underfitted the vector field, while spline interpolation introduced spurious local structures. RBF kernel interpolation performed most similarly to RVGP due to the homogeneous curvature of the human head. This observation is corroborated by significantly lower angular error and curl-divergence MAE for all subjects for RVGP compared with benchmarks ($n=61$, Fig. 3D). To test the variability against data density variation, we repeated the experiment for a 32-channel EEG signal and found that the errors were too large for all methods to be of practical relevance (Fig. S1). This shows that 64-channel EEG represents an empirical limit for resolving small-scale singularities.
@@
Given the superior reconstruction accuracy of RVGP, we asked if it could enhance the classification accuracy for patients with Alzheimer’s Disease (AD). Contemporary diagnostics for AD are costly, invasive, and laborious (Zetterberg & Bendlin, 2021). We instead employed a linear support vector machine to classify AD patients versus age-matched healthy controls based on the reconstructed divergence and curl fields derived from a brief, one-minute resting-state low-density EEG recording – a procedure that can be feasibly integrated into clinical settings due to its non-invasive nature and cost-efficiency. Our results show significantly higher than state-of-the-art classification accuracy, approaching that derived from the ground truth high-density EEG signal (Fig. 3E).
6 Discussion
@@
We introduced RVGP, a novel extension of Gaussian processes designed to model vector fields on latent Riemannian manifolds. Utilising the spectrum of the connection Laplacian operator, RVGP intrinsically captures the manifold’s geometry and topology and the vector field’s smoothness. This enables the method to learn global patterns while preserving singularities, filling a significant gap in existing approaches limited to known, analytically tractable manifolds. A key strength of RVGP is its data-driven, intrinsic construction via a proximity graph, which enhances its practical utility by making it highly applicable to real-world datasets where an explicit manifold parametrisation, e.g., for spheres and tori, is often unavailable. Demonstrated across diverse scientific domains such as neuroscience and geometric data analysis, RVGP advances the field of geometrically-informed probabilistic modelling and offers a statistical tool for various high-impact applications, including clinical neuroscience, that is robust to sampling density and noisy data.
As RVGP uses a proximity graph to approximate the manifold, one needs a sampling density that is high enough, combined with suitable similarity metric and graph algorithm to build an intrinsic approximation of the latent manifold, especially in high dimensions. Otherwise, RVGP will use the closest manifold approximated by thresholding the spectral decomposition of the connection Laplacian. Further, although we find that RVGP gives higher reconstruction accuracy than modelling the signal channel-wise on complex manifolds, as expected, this difference diminishes when the manifold curvature is homogeneous. In addition, we considered vector fields that lie in the tangent bundle of the manifold. However, one may also consider non-tangent but smooth vector fields by finding a (non-unique) $SO(d)$ rotation that maps the vectors into the tangent space, applying RVGP, and mapping back the predicted vectors by inverting the rotation. Future work may also study when the tangent spaces are not of uniform dimension by considering the sheaf Laplacian operator.
@@ A.1 Experimental data, and pre-processing
Experimental data: We analysed 10-minute resting state EEG recordings from a cohort of 61 subjects containing 33 patients with a clinical diagnosis of AD and 28 age-matched healthy control subjects. Data was acquired from the clinical imaging facility at Hammersmith Hospital in London. The AD group consisted of 15 Females and 18 Males with a mean age of 76 years (std = 7.8 years), healthy control group contained 15 Females and 13 Males with a mean age of 77 (std = 4.9 years). 10 participants were excluded from the original full cohort of 71 subjects due to poor data quality. During pre-processing, the EEG was bandpass filtered from 2-40 Hz with a second-order Butterworth filter and downsampled to 250Hz. Channels exceeding a kurtosis threshold of 3 were rejected before average re-referencing. We performed an independent component analysis with the Picard algorithm and applied an automatic artefact rejection algorithm in eeglab (Delorme & Makeig, 2004) with a maximum of 5 % of components removed from each dataset. The pre-processed EEG was then filtered in the alpha band (8-15Hz). The low-density EEG consisted of 64 channels sampled evenly across the scalp. These channels were then used to reconstruct the full 256-channel high-density recording.
@@ A.4 Feature Extraction
To determine the behaviourally relevant regions of the cortex, we computed the time-resolved divergence and vorticity fields for each subject. Next, we estimated the probability density of sources, sinks and spirals by computing the mean positive/negative vorticity and mean positive/negative divergence fields thresholded at $>1$ and $<-1$ averaged over time for each node. This produced four cortical probability maps per subject; mean source probability (positive divergence, $D>1$), mean sink probability (negative divergence $D<-1$), clockwise spiral probability (positive vorticity, $C>1$), anticlockwise spiral probability (negative vorticity, $C<-1$). We extracted two predictors from these four probability maps. The first predictor was the probability ratio of sources to sinks at a given node. The second predictor was the probability ratio of clockwise to anti-clockwise spirals at a given node:
@@
where $P_{n}(A)$ represents either the probability of a source or a clockwise spiral at node $n$ and $P_{n}(B)$ represents the probability of a sink or an anti-clockwise at node $n$. We ran a group-level cluster permutation analysis to identify nodes where the source-to-sink ratio or clockwise-to-anti-clockwise spiral ratio differed between the patients with AD and healthy controls in the reference EEG (alpha = 0.05, 50000 permutation). The three clusters with the lowest p-value were then used as seed regions where we extracted the source-to-sink probability ratio and clockwise to anti-clockwise spiral probability ratio in the interpolated EEG for each reconstruction method (RVGP, spline and linear interpolation).